# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ III-A4 CCC method
Observing (26), the initial phases of $\mathbf{S}^{S,b}$ on high and low-frequency bands are not aligned and there exists a different complex number $\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}$. To this end, a CCC method [23] is performed to eliminate the complex number $\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}$. Considering the noise, the procedure for applying the CCC method to $\mathbf{D}^{S,b}$ is presented in the following steps.
Step 1: $\mathbf{D}^{S,b}$ is divided into $N_{b}$ row vectors, denoted by
$\mathbf{D}^{S,b}=\left[\left(\mathbf{D}_{0}^{S,b}\right)^{\text{T}},\left(\mathbf{D}_{1}^{S,b}\right)^{\text{T}},\cdots,\left(\mathbf{D}_{N_{b}-1}^{S,b}\right)^{\text{T}}\right]^{\text{T}},$ (27)
where $\mathbf{D}_{n}^{S,b}\in\mathbb{C}^{1\times M_{b}}$ represents the $n$-th row vector of $\mathbf{D}^{S,b}$.
@@
Step 2: A delay feature vector $\mathbf{R}_{S}^{b}\in\mathbb{C}^{N_{b}\times 1}$ is obtained by accumulating the results of conjugate multiplication between the row vectors, denoted by (28), where $\mathbf{z}^{b}$ is an AWGN vector.
$\mathbf{R}_{S}^{b}$ $=\frac{1}{M_{b}}\left[\frac{1}{N_{b}}\sum_{a=0}^{N_{b}-1}\mathbf{D}_{a}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\frac{1}{N_{b}-1}\sum_{a=0}^{N_{b}-1-1}\mathbf{D}_{a+1}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\cdots,\frac{1}{N_{b}-n}\sum_{a=0}^{N_{b}-n-1}\mathbf{D}_{a+n}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\cdots,\mathbf{D}_{N_{b}-1}^{S,b}\left(\mathbf{D}_{0}^{S,b}\right)^{\text{H}}\right]^{\text{T}}$ (28)
$=\left[1,e^{-j2\pi\Delta f^{b}\tau_{0}},\cdots,e^{-j2\pi n\Delta f^{b}\tau_{0}},\cdots,e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}\right]^{\text{T}}+\mathbf{z}^{b}.$
@@
Step 3: $\mathbf{D}^{S,b}$ is divided into $M_{b}$ column vectors, expressed as
$\mathbf{D}^{S,b}=\left[\mathbf{C}_{0}^{S,b},\mathbf{C}_{1}^{S,b},\cdots,\mathbf{C}_{M_{b}-1}^{S,b}\right],$ (29)
@@
where $\mathbf{C}_{m}^{S,b}\in\mathbb{C}^{N_{b}\times 1}$ denotes the $m$-th column vector of $\mathbf{D}^{S,b}$.
@@
Step 4: A Doppler feature vector $\mathbf{V}_{S}^{b}\in\mathbb{C}^{1\times M_{b}}$ is obtained by accumulating the result of conjugate multiplication between the column vectors, expressed as (30).
$\mathbf{V}_{S}^{b}$ $=\frac{1}{N_{b}}\left[\frac{1}{M_{b}}\sum_{a=0}^{M_{b}-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a}^{S,b},\frac{1}{M_{b}-1}\sum_{a=0}^{M_{b}-1-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a+1}^{S,b},\cdots,\frac{1}{M_{b}-n}\sum_{a=0}^{M_{b}-n-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a+n}^{S,b},\cdots,\left(\mathbf{C}_{0}^{S,b}\right)^{\text{H}}\mathbf{C}_{N_{b}-1}^{S,b}\right]$ (30)
$=\left[1,e^{j2\pi\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}},\cdots,e^{j2\pi m\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}},\cdots,e^{j2\pi(M_{b}-1)\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}}\right]+\mathbf{\hat{z}}^{b}.$
After the above steps, the four feature vectors are obtained, namely $\mathbf{R}_{S}^{1}$, $\mathbf{V}_{S}^{1}$, $\mathbf{R}_{S}^{2}$, and $\mathbf{V}_{S}^{2}$.
@@ III-B Sensing Information Fusion Stage
In the sensing information fusion stage, we fuse the feature vectors obtained in the signal preprocessing stage by the symbol-level fusion and estimate the parameters of target, which include the fusion algorithm of delay feature vectors and the fusion algorithm of Doppler feature vectors. Without loss of generality, the following remarks need to be specified.
@@ III-B1 Fusion algorithm of delay feature vectors
Without considering the noise, observing (28), the difference between $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$ is the subcarrier spacing, which makes the symbol-level fusion difficult. According to $\frac{\Delta f^{2}}{\Delta f^{1}}=Q$, $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$ are rewritten as
$\mathbf{R}_{S}^{1}$ $=\left[e^{-j2\pi n\Delta f^{1}\tau_{0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{1}-1},$ (31)
$\mathbf{R}_{S}^{2}$ $=\left[e^{-j2\pi nQ\Delta f^{1}\tau_{0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{2}-1}.$ (32)
@@
$\hat{\mathbf{R}}_{S}^{2}=\left[e^{-j2\pi n^{\prime\prime}\Delta f^{1}\tau_{0}}\right]^{\text{T}}\in\mathbb{C}^{N_{2}\times 1}|_{n^{\prime\prime}=0,Q,\cdots,Q(N_{2}-1)}.$ (33)
Then, the subcarrier spacings in (31) and (33) are identical. The fusion process of delay feature vectors is discussed below in two cases.
Case 1 $(N_{1}-1<Q(N_{2}-1))$: Initialize a zero-element vector $\mathbf{P}_{S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$ and an index value $\xi_{1}\in\{0,1,\cdots,QN_{2}-1\}$. Then, we traverse the element of $\mathbf{P}_{S}^{1}$ with $\xi_{1}$. When $\xi_{1}$ equals $n$, the element $\mathbf{R}_{S}^{1}(n)$ is assigned to $\mathbf{P}_{S}^{1}(\xi_{1})$. When $\xi_{1}$ equals $n^{\prime\prime}$, the element $\hat{\mathbf{R}}_{S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{P}_{S}^{1}(\xi_{1})$. Given that the zero elements in $\mathbf{P}_{S}^{1}$ deteriorate the Fourier sidelobes, which can be mitigated by recovering the missing sensing data with the CS-based sensing method proposed in [11]. The final processed vector is denoted by $\hat{\mathbf{P}}_{S}^{1}$.
Case 2 $\left(N_{1}-1\geq Q(N_{2}-1)\right)$: Initialize a zero-element vector $\mathbf{P}_{S}^{2}\in\mathbb{C}^{N_{1}\times 1}$ and an index value $\xi_{2}\in\{0,1,\cdots,N_{1}-1\}$. Then, we traverse the element of $\mathbf{P}_{S}^{2}$ with $\xi_{2}$. When $\xi_{2}$ equals $n$, the element $\mathbf{R}_{S}^{1}(n)$ is assigned to $\mathbf{P}_{S}^{2}(\xi_{2})$. When $\xi_{2}$ equals $n^{\prime\prime}$, the element $\hat{\mathbf{R}}_{S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{P}_{S}^{2}(\xi_{2})$.
Taking the $\hat{\mathbf{P}}_{S}^{1}$ in Case 1 as an example, an improved IDFT method is used to achieve a high-accuracy sensing. The specific procedures are as follows, while the Case 2 has the similar procedures.
@@
Step 2: Generate a distance searching vector $\vec{\mathbf{A}}$ by gridding the searching interval with a size of grid being $\Delta R=\frac{R_{\text{max}}-R_{\text{min}}}{J}$, where $J$ is the number of girds and $\vec{\mathbf{A}}\in\mathbb{C}^{J\times 1}$ is denoted by
$\vec{\mathbf{A}}=\left[R_{1},R_{2},\cdots,R_{\eta},\cdots,R_{J}\right]^{\text{T}},$ (34)
@@
Step 3: According to the size of $\hat{\mathbf{P}}_{S}^{1}$ and the expression form of delay feature vectors in (31), the $\vec{\mathbf{A}}$ is transformed to a searching matrix $\mathbf{A}_{1}\in\mathbb{C}^{J\times QN_{2}}$, which is expressed as
@@
Step 4: Obtain and search the delay profile $\vec{\mathbf{P}}_{1}=\mathbf{A}_{1}\hat{\mathbf{P}}_{S}^{1}$. The peak index of $\vec{\mathbf{P}}_{1}$ is denoted by $\hat{\eta}_{1}$.
Step 5: The range estimation of target is $\hat{r}_{0}^{1}=\vec{\mathbf{A}}(\hat{\eta}_{1})$.
The fusion algorithm of delay feature vectors is shown in Algorithm 1, which is explained intuitively in Fig. 3.
@@
Delay feature vectors $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$;
@@
Output: The range estimations of target $\hat{r}_{0}^{1}$ and $\hat{r}_{0}^{2}$.
@@
2:     Assume a vector $\mathbf{P}_{S}^{2}\in\mathbb{C}^{N_{1}\times 1}=\mathbf{R}_{S}^{1}$,
@@
3:     for $\mathbf{P}_{S}^{2}$ $\mu_{2}$ in $\mathcal{N}_{2}$ do
4:        The element value of $\mathbf{R}_{S}^{2}(\mu_{2})$ is added to $\mathbf{P}_{S}^{2}(Q\mu_{2})$;
5:        $\mathbf{P}_{S}^{2}(Q\mu_{2})$ = $\mathbf{P}_{S}^{2}(Q\mu_{2})/2$;
@@
8:     Initialize a zero-element vector $\mathbf{P}_{S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$,
@@
9:     for $\mathbf{P}_{S}^{1}$ $\mu_{1}$ in $\mathcal{N}_{1}$ do
10:        The element value of $\mathbf{R}_{S}^{1}(\mu_{1})$ is assigned to $\mathbf{P}_{S}^{1}(\mu_{1})$;
@@
13:     for $\mathbf{P}_{S}^{1}$ $\mu_{3}$ in $\mathcal{N}_{3}$ do
14:        The element value of $\mathbf{R}_{S}^{2}(\mu_{3})$ is added to $\mathbf{P}_{S}^{1}(Q\mu_{3})$;
@@
16:           $\mathbf{P}_{S}^{1}(Q\mu_{3})$ = $\mathbf{P}_{S}^{1}(Q\mu_{3})/2$;
@@
19:     Recover the missing elements of $\mathbf{P}_{S}^{1}$ by
    CS-based method proposed in [11] to obtain $\hat{\mathbf{P}}_{S}^{1}$;
@@
21: The results of $\mathbf{P}_{S}^{1}$ and $\mathbf{P}_{S}^{2}$ undergoing improved IDFT are
searched for peaks to obtain the peak indices $\hat{\eta}^{1}$ and $\hat{\eta}^{2}$;
22: Substitute $\hat{\eta}^{1}$ and $\hat{\eta}^{2}$ into (34) to obtain the range
estimations of target $\hat{r}_{0}^{1}$ and $\hat{r}_{0}^{2}$.

### v2

@@ III-A3 CCC method
For the $i^{\prime}$-th target, observing (21), the initial phases of $\mathbf{S}_{i^{\prime}}^{S,b}$ on high and low-frequency bands are not aligned and there exists a different complex number $\kappa_{i^{\prime},S}^{b}$. To this end, a CCC method [18] is performed to eliminate the complex number $\kappa_{i^{\prime},S}^{b}$. The procedure for applying the CCC method to $\mathbf{D}_{i^{\prime}}^{S,b}$ is presented in the following steps.
Step 1: $\mathbf{D}_{i^{\prime}}^{S,b}$ is divided into $N_{b}$ row vectors, denoted by
$\mathbf{D}_{i^{\prime}}^{S,b}=\left[\left(\mathbf{d}_{i^{\prime},0}^{S,b}\right)^{\text{T}},\left(\mathbf{d}_{i^{\prime},1}^{S,b}\right)^{\text{T}},\cdots,\left(\mathbf{d}_{i^{\prime},N_{b}-1}^{S,b}\right)^{\text{T}}\right]^{\text{T}},$ (22)
where $\mathbf{d}_{i^{\prime},n}^{S,b}\in\mathbb{C}^{1\times M_{b}}$ represents the $n$-th row vector of $\mathbf{D}_{i^{\prime}}^{S,b}$.
@@
Step 2: A delay feature vector $\mathbf{r}_{i^{\prime},S}^{b}\in\mathbb{C}^{N_{b}\times 1}$ is obtained by accumulating the results of conjugate multiplication between the row vectors, denoted by (23), where $\mathbf{z}_{i^{\prime}}^{b}$ is an interference vector.
$\mathbf{r}_{i^{\prime},S}^{b}$ $=\frac{1}{M_{b}}\left[\frac{1}{N_{b}}\sum_{a=0}^{N_{b}-1}\mathbf{d}_{i^{\prime},a}^{S,b}\left(\mathbf{d}_{i^{\prime},a}^{S,b}\right)^{\text{H}},\frac{1}{N_{b}-1}\sum_{a=0}^{N_{b}-1-1}\mathbf{d}_{i^{\prime},a+1}^{S,b}\left(\mathbf{d}_{i^{\prime},a}^{S,b}\right)^{\text{H}},\cdots,\frac{1}{N_{b}-n}\sum_{a=0}^{N_{b}-n-1}\mathbf{d}_{i^{\prime},a+n}^{S,b}\left(\mathbf{d}_{i^{\prime},a}^{S,b}\right)^{\text{H}},\cdots,\mathbf{d}_{i^{\prime},N_{b}-1}^{S,b}\left(\mathbf{d}_{i^{\prime},0}^{S,b}\right)^{\text{H}}\right]^{\text{T}}$ (23)
$=\left[1,e^{-j2\pi\Delta f^{b}\tau_{i^{\prime},0}},\cdots,e^{-j2\pi n\Delta f^{b}\tau_{i^{\prime},0}},\cdots,e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{i^{\prime},0}}\right]^{\text{T}}+\mathbf{z}_{i^{\prime}}^{b}.$
To guarantee that the SNR of $\mathbf{r}_{i^{\prime},S}^{b}$ is evenly distributed, weighting averaging is needed to obtain the final feature vector $\mathbf{g}_{i^{\prime},S}^{b}=\mathbf{r}_{i^{\prime},S}^{b}\odot\mathbf{w}_{r}$, where $\mathbf{w}_{r}=[\frac{N_{b}}{N_{b}(N_{b}+1)/2},\frac{N_{b}-1}{N_{b}(N_{b}+1)/2},\cdots,\frac{1}{N_{b}(N_{b}+1)/2}]^{\text{T}}$ is the weight vector.
@@
Step 3: $\mathbf{D}_{i^{\prime}}^{S,b}$ is divided into $M_{b}$ column vectors, expressed as
$\mathbf{D}_{i^{\prime}}^{S,b}=\left[\mathbf{c}_{i^{\prime},0}^{S,b},\mathbf{c}_{i^{\prime},1}^{S,b},\cdots,\mathbf{c}_{i^{\prime},M_{b}-1}^{S,b}\right],$ (24)
@@
where $\mathbf{c}_{i^{\prime},m}^{S,b}\in\mathbb{C}^{N_{b}\times 1}$ denotes the $m$-th column vector of $\mathbf{D}_{i^{\prime}}^{S,b}$.
@@
Step 4: A Doppler feature vector $\mathbf{v}_{i^{\prime},S}^{b}\in\mathbb{C}^{1\times M_{b}}$ is obtained by accumulating the result of conjugate multiplication between the column vectors, expressed as (25).
$\mathbf{v}_{i^{\prime},S}^{b}$ $=\frac{1}{N_{b}}\left[\frac{1}{M_{b}}\sum_{a=0}^{M_{b}-1}\left(\mathbf{c}_{i^{\prime},a}^{S,b}\right)^{\text{H}}\mathbf{c}_{i^{\prime},a}^{S,b},\frac{1}{M_{b}-1}\sum_{a=0}^{M_{b}-1-1}\left(\mathbf{c}_{i^{\prime},a}^{S,b}\right)^{\text{H}}\mathbf{c}_{i^{\prime},a+1}^{S,b},\cdots,\frac{1}{M_{b}-n}\sum_{a=0}^{M_{b}-n-1}\left(\mathbf{c}_{i^{\prime},a}^{S,b}\right)^{\text{H}}\mathbf{c}_{i^{\prime},a+n}^{S,b},\cdots,\left(\mathbf{c}_{i^{\prime},0}^{S,b}\right)^{\text{H}}\mathbf{c}_{i^{\prime},N_{b}-1}^{S,b}\right]$ (25)
$=\left[1,e^{j2\pi\left(\frac{2f_{\mathrm{C}}^{b}v_{i^{\prime},0}}{c}\right)T^{b}},\cdots,e^{j2\pi m\left(\frac{2f_{\mathrm{C}}^{b}v_{i^{\prime},0}}{c}\right)T^{b}},\cdots,e^{j2\pi(M_{b}-1)\left(\frac{2f_{\mathrm{C}}^{b}v_{i^{\prime},0}}{c}\right)T^{b}}\right]+\mathbf{\hat{z}}_{i^{\prime}}^{b}.$
Similar to delay feature vector, the final Doppler vector is $\mathbf{e}_{i^{\prime},S}^{b}=\mathbf{v}_{i^{\prime},S}^{b}\odot\mathbf{w}_{v}$, where $\mathbf{w}_{v}=[\frac{M_{b}}{M_{b}(M_{b}+1)/2},\frac{M_{b}-1}{M_{b}(M_{b}+1)/2},\cdots,\frac{1}{M_{b}(M_{b}+1)/2}]$ is the weight vector. After the above steps, the feature vector sets of the $I$ targets are obtained, namely $\{\mathbf{g}_{i^{\prime},S}^{1}\}_{i^{\prime}=1,2,\cdots,I}$, $\{\mathbf{e}_{i^{\prime},S}^{1}\}_{i^{\prime}=1,2,\cdots,I}$, $\{\mathbf{g}_{i^{\prime},S}^{2}\}_{i^{\prime}=1,2,\cdots,I}$, and $\{\mathbf{e}_{i^{\prime},S}^{2}\}_{i^{\prime}=1,2,\cdots,I}$.
@@ III-B Sensing Information Fusion Stage
In the sensing information fusion stage, we fuse the feature vectors obtained in the signal preprocessing stage by the symbol-level fusion and estimate the parameters of $I$ targets, which include the fusion algorithm of delay feature vectors and the fusion algorithm of Doppler feature vectors. Without loss of generality, the following remarks need to be specified.
@@ III-B1 Fusion algorithm of delay feature vectors
Taking the $\mathbf{g}_{i^{\prime},S}^{1}$ and $\mathbf{g}_{i^{\prime},S}^{2}$ of the $i^{\prime}$-th target as an example. Without considering the interference, observing (23), the difference between $\mathbf{g}_{i^{\prime},S}^{1}$ and $\mathbf{g}_{i^{\prime},S}^{2}$ is the subcarrier spacing, which makes the symbol-level fusion difficult. According to $\frac{\Delta f^{2}}{\Delta f^{1}}=Q$, $\mathbf{g}_{i^{\prime},S}^{1}$ and $\mathbf{g}_{i^{\prime},S}^{2}$ are rewritten as
$\mathbf{g}_{i^{\prime},S}^{1}$ $=\left[e^{-j2\pi n\Delta f^{1}\tau_{i^{\prime},0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{1}-1},$ (26)
$\mathbf{g}_{i^{\prime},S}^{2}$ $=\left[e^{-j2\pi nQ\Delta f^{1}\tau_{i^{\prime},0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{2}-1}.$ (27)
@@
$\hat{\mathbf{g}}_{i^{\prime},S}^{2}=\left[e^{-j2\pi n^{\prime\prime}\Delta f^{1}\tau_{i^{\prime},0}}\right]^{\text{T}}\in\mathbb{C}^{N_{2}\times 1}|_{n^{\prime\prime}=0,Q,\cdots,Q(N_{2}-1)}.$ (28)
Then, the subcarrier spacings in (26) and (28) are identical. According to the MRC principle, the delay feature vectors are multiplied by the weight to carry out the subsequent symbol-level fusion, and $\tilde{\mathbf{r}}_{i^{\prime},S}^{1}=\frac{\sigma_{S,2}^{2}}{\sigma_{S,1}^{2}+\sigma_{S,2}^{2}}\mathbf{g}_{i^{\prime},S}^{1}$ and $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}=\frac{\sigma_{S,1}^{2}}{\sigma_{S,1}^{2}+\sigma_{S,2}^{2}}\hat{\mathbf{g}}_{i^{\prime},S}^{2}$ are obtained. The fusion process of delay feature vectors is discussed below in two cases.
Case 1 $(N_{1}-1<Q(N_{2}-1))$: Initialize a zero-element vector $\mathbf{p}_{i^{\prime},S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$ and an index value $\xi_{1}\in\{0,1,\cdots,QN_{2}-1\}$. Then, we traverse the element of $\mathbf{p}_{i^{\prime},S}^{1}$ with $\xi_{1}$. When $\xi_{1}$ equals $n$, the element $\tilde{\mathbf{r}}_{i^{\prime},S}^{1}(n)$ is assigned to $\mathbf{p}_{i^{\prime},S}^{1}(\xi_{1})$. When $\xi_{1}$ equals $n^{\prime\prime}$, the element $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{p}_{i^{\prime},S}^{1}(\xi_{1})$. Given that the zero elements in $\mathbf{p}_{i^{\prime},S}^{1}$ deteriorate the Fourier sidelobes, which can be mitigated by recovering the missing sensing data with the CS-based sensing method proposed in [15]. The final processed vector is denoted by $\hat{\mathbf{p}}_{i^{\prime},S}^{1}$.
Case 2 $\left(N_{1}-1\geq Q(N_{2}-1)\right)$: Initialize a zero-element vector $\mathbf{p}_{i^{\prime},S}^{2}\in\mathbb{C}^{N_{1}\times 1}$ and an index value $\xi_{2}\in\{0,1,\cdots,N_{1}-1\}$. Then, we traverse the element of $\mathbf{p}_{i^{\prime},S}^{2}$ with $\xi_{2}$. When $\xi_{2}$ equals $n$, the element $\tilde{\mathbf{r}}_{i^{\prime},S}^{1}(n)$ is assigned to $\mathbf{p}_{i^{\prime},S}^{2}(\xi_{2})$. When $\xi_{2}$ equals $n^{\prime\prime}$, the element $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{p}_{i^{\prime},S}^{2}(\xi_{2})$.
Taking the $\hat{\mathbf{p}}_{i^{\prime},S}^{1}$ in Case 1 as an example, an improved IDFT method is used to achieve a high-accuracy sensing. The specific procedures are as follows, while the Case 2 has the similar procedures.
@@
Step 2: Generate a distance searching vector $\mathbf{a}$ by gridding the searching interval with a size of grid being $\Delta R=\frac{R_{\text{max}}-R_{\text{min}}}{J}$, where $J$ is the number of girds and $\mathbf{a}\in\mathbb{C}^{J\times 1}$ is denoted by
$\mathbf{a}=\left[R_{1},R_{2},\cdots,R_{\eta},\cdots,R_{J}\right]^{\text{T}},$ (29)
@@
Step 3: According to the size of $\hat{\mathbf{p}}_{i^{\prime},S}^{1}$ and the expression form of delay feature vectors in (26), the $\mathbf{a}$ is transformed to a searching matrix $\mathbf{A}_{1}\in\mathbb{C}^{J\times QN_{2}}$, which is expressed as
@@
Step 4: Obtain and search the delay profile $\tilde{\mathbf{p}}_{i^{\prime},1}=\mathbf{A}_{1}\hat{\mathbf{p}}_{i^{\prime},S}^{1}$. The peak index of $\tilde{\mathbf{p}}_{i^{\prime},1}$ is denoted by $\hat{\eta}_{i^{\prime}}^{1}$.
Step 5: The range estimation of the $i^{\prime}$-th target is $\hat{r}_{i^{\prime},0}^{1}=\mathbf{a}(\hat{\eta}_{i^{\prime}}^{1})$.
The fusion algorithm of delay feature vectors for the $i$-th target is shown in Algorithm 1, which is explained intuitively in Fig. 3. Finally, the estimated ranges of $I$ target are obtained.
@@
Delay feature vectors of the $i^{\prime}$-th target $\tilde{\mathbf{r}}_{i^{\prime},S}^{1}$ and $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}$;
@@
Output: The range estimations of the $i^{\prime}$-th target $\hat{r}_{i^{\prime},0}^{1}$ and $\hat{r}_{i^{\prime},0}^{2}$.
@@
2:     Assume a vector $\mathbf{p}_{i^{\prime},S}^{2}\in\mathbb{C}^{N_{1}\times 1}=\tilde{\mathbf{r}}_{i^{\prime},S}^{1}$,
@@
3:     for $\mu_{2}$ in $\mathcal{N}_{2}$ do
4:        The element value of $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}(\mu_{2})$ is added to $\mathbf{p}_{i^{\prime},S}^{2}(Q\mu_{2})$;
5:        $\mathbf{p}_{i^{\prime},S}^{2}(Q\mu_{2})$ =$\mathbf{p}_{i^{\prime},S}^{2}(Q\mu_{2})/2$;
@@
8:     Initialize a zero-element vector $\mathbf{p}_{i^{\prime},S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$,
@@
9:     for $\mu_{1}$ in $\mathcal{N}_{1}$ do
10:        The element value of $\tilde{\mathbf{r}}_{i^{\prime},S}^{1}(\mu_{1})$ is assigned to $\mathbf{p}_{i^{\prime},S}^{1}(\mu_{1})$;
@@
13:     for $\mu_{3}$ in $\mathcal{N}_{3}$ do
14:        The element value of $\tilde{\mathbf{r}}_{i^{\prime},S}^{2}(\mu_{3})$ is added to $\mathbf{p}_{i^{\prime},S}^{1}(Q\mu_{3})$;
@@
16:           $\mathbf{p}_{i^{\prime},S}^{1}(Q\mu_{3})$ =$\mathbf{p}_{i^{\prime},S}^{1}(Q\mu_{3})/2$;
@@
19:     Recover the missing elements of $\mathbf{p}_{i^{\prime},S}^{1}$ by
    CS-based method proposed in [15] to obtain $\hat{\mathbf{p}}_{i^{\prime},S}^{1}$;
@@
21: The results of $\hat{\mathbf{p}}_{i^{\prime},S}^{1}$ and $\mathbf{p}_{i^{\prime},S}^{2}$ undergoing improved IDFT are
searched for peaks to obtain the peak indices $\hat{\eta}_{i^{\prime}}^{1}$ and $\hat{\eta}_{i^{\prime}}^{2}$;
22: Substitute $\hat{\eta}_{i^{\prime}}^{1}$ and $\hat{\eta}_{i^{\prime}}^{2}$ into (29) to obtain the range
estimations of the $i^{\prime}$-th target $\hat{r}_{i^{\prime},0}^{1}$ and $\hat{r}_{i^{\prime},0}^{2}$.